### v2

# MTFM: A Scalable and Alignment-free Foundation Model for Industrial Recommendation in Meituan

Abstract.
Industrial recommendation systems typically involve multiple scenarios, yet existing cross-domain (CDR) and multi-scenario (MSR) methods often require prohibitive resources and strict input alignment, limiting their extensibility. We propose MTFM (Meituan Foundation Model for Recommendation), a transformer-based framework that addresses these challenges. Instead of pre-aligning inputs, MTFM transforms cross-domain data into heterogeneous tokens, capturing multi-scenario knowledge in an alignment-free manner. To enhance efficiency, we first introduce a multi-scenario user-level sample aggregation that significantly enhances training throughput by reducing the total number of instances. We further integrate Grouped-Query Attention and a customized Hybrid Target Attention to minimize memory usage and computational complexity. Furthermore, we implement various system-level optimizations, such as kernel fusion and the elimination of CPU-GPU blocking, to further enhance both training and inference throughput. Offline and online experiments validate the effectiveness of MTFM, demonstrating that significant performance gains are achieved by scaling both model capacity and multi-scenario training data.
1. Introduction
The landscape of Large Language Models (LLM) has been redefined by the era of Foundation Models (FMs), where unified architectures have evolved from unimodal to multimodal (Bai et al., 2025; Yang et al., 2025; Team et al., 2025; Fu et al., 2025; Liu et al., 2026). In parallel, the recommendation community has begun to validate the Scaling Laws (Kaplan et al., 2020; Hoffmann et al., 2022), primarily focusing on the expansion of model parameters and data volume within isolated scenarios (Zhai et al., 2024; Zhu et al., 2025; Zhang et al., 2024b; Wang et al., 2025; Zhou et al., 2025a; Han et al., 2024). However, the true potential of a Recommendation Foundation Model remains untapped as long as models are confined to single-scenario silos. Inspired by the success of multimodal FMs, we argue that the next frontier lies in bridging cross-scenario heterogeneity. By extending from single-scene optimization to a unified multi-scenario framework, we can leverage diverse behavioral signals to further unlock the benefits of Scaling Laws, thereby elevating the performance upper bound and establishing a more scalable paradigm for recommendation.
To realize this vision, we contend that an ideal Recommendation Foundation Model must embody three quintessential properties: Scalability, Extensibility, and Efficiency. First, Scalability requires the model to yield consistent and predictable performance gains as the parameter size and data volume expand. Second, Extensibility dictates that the model can seamlessly adapt to an arbitrary number of existing scenarios or integrate emerging ones with low cost. Finally, Efficiency; the framework must maintain low computational overhead during both training and inference, even when handling the massive data volumes generated across diverse scenarios.
Existing multi-scenario modeling efforts (Sheng et al., 2021a; Shu et al., 2024; Zhang et al., 2024a; Xiaoyu et al., 2025) primarily follow a ”harmonize-then-decompose” paradigm. These approaches typically aggregate data from multiple scenarios and enforce input homogeneity with a fixed template. Any heterogeneous components must be either discarded or handled through an additional padding operator. Structurally, they rely on disentangling model parameters into domain-invariant and domain-specific components (e.g., via MoE-based architectures (Ma et al., 2018a; Fedus et al., 2022)) to capture shared knowledge while preserving scenario-specific nuances.
However, these conventional paradigms fall short of the aforementioned hallmarks for a foundation model. 1) Rigidity in Extensibility: Industrial recommendations often encompass highly divergent scenarios with heterogeneous feature spaces. For instance, in our food delivery platform, the feature schemas for restaurant recommendation and food recommendations are inherently distinct due to different types of supply and UI presentations. The reliance on strict feature alignment is unsustainable in industrial ecosystems. Forcing hundreds of heterogeneous features into a fixed template is not only error-prone but also leads to significant information loss. 2) Lack of Architectural Scalability: Most existing structures are handcrafted based on expert heuristics to handle a fixed set of scenarios. Such intricate, scenario-tailored designs lack the structural flexibility needed to benefit from the scaling laws observed in other scenarios. 3) Prohibitive Computational Costs: Within the traditional paradigm, training costs increase linearly with data volume. The brute-force integration of massive multi-scenario data leads to exorbitant expenses, rendering the model economically inefficient for real-world deployment.
To address these challenges, we present the Meituan Foundation Model (MTFM), a unified framework engineered to satisfy the three quintessential properties. First, to ensure Extensibility, MTFM moves beyond rigid feature templates by abstracting multi-domain inputs into a unified sequence of heterogeneous tokens. This token-based representation allows the model to ingest disparate signals across scenarios without manual alignment or information loss. Building upon this, we employ a transformer-inspired backbone designed for Scalability. By replacing handcrafted heuristic-based structures with a deep self-attention mechanism, MTFM empowers the model to automatically capture both universal behavioral laws and scenario-specific nuances as the model and data scale.
To achieve superior Efficiency without compromising modeling capacity, we introduce Hybrid Target Attention (HTA). While full-attention backbones excel in expressiveness, their quadratic complexity is prohibitive for industrial-scale long sequences. Inspired by the application of sparse attention in LLMs (Yuan et al., 2025; Qwen Team, 2025; Team et al., 2026), HTA serves as a strategic middle ground: it selectively applies full attention to a few critical layers to maintain global dependency capture, while utilizing target attention for the majority of the stack. By further integrating Grouped-Query Attention (GQA) (Ainslie et al., 2023), MTFM significantly compresses the memory footprint and improves the training throughput, enabling the processing of more multi-scenario tokens with minimal overhead.
Beyond architectural innovations, we implement a suite of systematic optimizations to bridge the gap between theoretical models and real-world deployment. On the data front, we utilize user-level aggregation from multiple scenarios to transform fragmented samples into dense sequences, substantially boosting training throughput. At the system level, we nearly double both training and inference throughput by eliminating critical bottlenecks through a range of techniques, such as CPU-GPU pipeline orchestration, kernel fusion, and structured sparsity. These co-designs ensure that MTFM is economically viable for large-scale production.
We evaluate the performance of MTFM through both offline experiments and online A/B testing across various scenarios in Meituan. Offline experiments demonstrate that MTFM consistently outperforms state-of-the-art baselines across multiple scenarios and objectives. Specifically, for the CTR (Click-Through Rate) task, MTFM achieves an average GAUC (Grouped Area Under the Curve) improvement of 0.36pp (percentage points), with a maximum gain of 0.76pp. For CTCVR (Click-Through Conversion Rate) tasks, the model yields an average GAUC increase of 0.29pp and a peak improvement of 0.53pp. Furthermore, Real-world online A/B testing shows significant business growth: a +2.98% increase in orders for the Shenqiangshou (SQS) Coupon-Package Recommendation, and a +1.45% boost for the Pinhaofan (PHF) Food Recommendation.
Our main contributions are summarized as follows:
Unified Foundation Architecture: We propose MTFM, a multi-scenario foundation model that uses heterogeneous tokenization to ensure seamless extensibility and scalability across disparate domains.
Hybrid Target Attention: We introduce HTA, a novel attention mechanism combining sparse-dense layers and GQA to achieve an optimal trade-off between modeling capacity and computational efficiency.
System-Model Co-design: We implement system-level optimizations to ensure the economic viability of MTFM in industrial-scale production.
Empirical Validation: Extensive offline and online evaluations at Meituan demonstrate significant gains, proving MTFM’s real-world effectiveness.
2. Related Work
2.1. Multi-Scenario Recommendation
Existing research primarily follows a ”harmonize-then-decompose” paradigm, which aims to balance commonality and diversity by disentangling model parameters into domain-invariant and domain-specific components. For instance, STAR (Sheng et al., 2021a) pioneered this approach with a star topology that anchors domain-specific parameters around shared centered weights. To further capture complex feature interactions, M3OE (Zhang et al., 2024a) leverages three specialized MoE modules to learn hierarchical user preferences across common, domain, and task aspects. More recently, MLoRA (Yang et al., 2024) introduces a more parameter-efficient alternative by adapting LoRA modules for individual domains.
Moving beyond parameter-level disentanglement, recent studies have increasingly explored multi-scenario foundation models to achieve superior generalization. This line of work often adopts a ”Foundation-Expert” architecture. For example, ExFM (Recommendation, 2025) distills knowledge from ultra-large-scale models into scenario-specific experts, while LFM4Ads (Zhang et al., 2025b) utilizes multi-granularity mechanisms to transfer comprehensive representations across features and tasks. Similarly, (Li et al., 2025) incorporates HSTU as a robust encoder to bridge foundational target information with downstream expert models. However, these approaches predominantly rely on a two-stage pipeline rather than a fully end-to-end foundation model, which potentially limits the model’s performance upper bound.
2.2. Ranking Model with Scalability
Current methods are primarily categorized into two paradigms: Transformer-based and non-Transformer-based (often referred to as Deep Learning Recommendation Models, DLRM). Transformer-based approaches partition input features into individual tokens and leverage Transformer-like architectures for unified modeling. Specifically, HSTU (Zhai et al., 2024) reconceptualizes recommendation as a sequential transduction task, employing a high-performance architecture to capture user behavior patterns. To further enrich feature interactions, MTGR (Han et al., 2025) extends this by incorporating cross-features and bidirectional attention mechanisms. Recognizing the computational demands of such models, OneTrans (Zhang et al., 2025a) addresses inference efficiency via a pyramid-style tail truncation strategy. More recently, scaling laws have been explored within generative frameworks (Deng et al., 2025; Zhou et al., 2025b), and these efforts also remain rooted in the Transformer backbone. Conversely, traditional DLRM adopts a modular design, decomposing the system into distinct functional components. These methods achieve scalability by selectively expanding specific stackable modules. For instance, RankMixer (Zhu et al., 2025) enhances the classic DLRM architecture by integrating Per-token Feed-Forward Networks to boost model capacity. Similarly, WuKong (Zhang et al., 2024b) attains scalability through a specialized design of stacked layers, combining Factorization Machine blocks with Linear Compress blocks to effectively increase model scalability.
3. Methodology
In this section, we first introduce the multi-scenario data arrangement method in Section 3.1. Next, we elaborate on the MTFM model architecture in Section 3.2, including the heterogeneous tokenization module and hybrid attention architecture. Finally, we discuss optimization techniques during training and deployment in Section 3.3, which are critical for successfully applying MTFM to industrial-scale multi-scenario recommendation systems.
3.1. User Sample Aggregation with Multi-Scenario Data
In this subsection, we extend the data arrangement method from MTGR (Han et al., 2025) for multi-scenario modeling. Following MTGR, we aggregate samples at the user level during training and at the request level during inference. The difference is that in MTFM, the aggregated samples contain candidates and features from multiple scenarios.
For offline training, we aggregate all exposure behaviors in a specific time window from multiple scenarios at the user-level. Specifically, each user corresponds to multiple training samples, denoted by , where and are the sets of input features and labels, respectively. In the following, we omit the subscript for brevity when there is no ambiguity. In more detail, the input feature set and the label set with user-level aggregation are denoted as follows:
where , and are the sets of cross features, item-side features, and labels of the -th exposure in scenario . is the number of exposures in scenario , and is the number of scenarios. is the set of scenario-specific user features in scenario , is the set of historical interaction sequences, and records the user’s most recent behaviors. Each or represents an action, including multiple features such as item ID, item tag, etc., and is the sequence length. Since the timing of behaviors in may overlap with the sample aggregation window, some actions might occur after the exposure represented by and . Modeling these directly introduces the risk of information leakage. Following the approach in MTGR, we will address this by constructing a dynamic mask based on interaction timestamps in the next section.
Notably, as user behavior sequences across diverse scenarios constitute integral parts of their overall interests, both historical and real-time sequences ( and ) are shared across all scenarios. In contrast, for scenario-specific features—including user attributes , cross-features , and item-side characteristics are treated as scenario-specific. This configuration is dictated by the fundamental structural divergence between domains; for instance, the restaurant recommendation scenario naturally lacks specific food tags, rendering the feature schemas across these scenarios intrinsically incompatible for direct alignment.
For online inference, each scenario deploys its corresponding subgraph in MTFM and aggregates the features of all candidates at the request level. Our subgraph restricts computation to components relevant to the current scenario, bypassing scenario-specific modules belonging to other scenarios. Specically, the input feature for scenario is denoted by , where is the number of candidates in one request of scenario .
3.2. MTFM Model Architecture
In this subsection, we introduce the model architecture of MTFM as shown in Figure 1.
3.2.1. Heterogeneous Tokenization
We first detail how our model represents the input features as a variable-length sequence of heterogeneous tokens.
There are three primary types of tokens in our model, including H-token, R-token, and T-token. Taking as an example, each item in corresponds to an H-token, whose embedding is generated by first embedding raw item features and subsequently applying an MLP to project them into a unified dimension :
Different are employed as tokenizers for the H-tokens corresponding to different historical sequences, due to the heterogeneity of features contained within these sequences.
After that, we sort the token embeddings of items in all historical sequences in chronological order to form a new embedding matrix , where is the total number of items in all historical sequences.
Similarly, the real-time sequences are tokenized into an embedding matrix , where is the total number of items in all real-time sequences.
Each exposure behavior corresponds to a T-token, whose input features are the concatenation of the user profile, cross and target item features. Specically, the token embedding of an exposure behavior on a target item is obtained as follows:
where means column concatenation. All the exposure behaviors are tokenized into an embedding matrix , where is the total number of exposures in all scenarios.
Finally, all the H-, R-, and T-tokens are stacked to form the initial input embedding matrix to our Hybrid Attention Architecture:
where means row concatenation and is the total number of tokens which can vary for different users.
3.2.2. Hybrid Target Attention Architecture
While Transformers are natural candidates for modeling the resulting heterogeneous token sequences, their complexity poses significant computational challenges in multi-scenario settings where sequence lengths increase substantially. To balance modeling efficacy and efficiency, we propose a Hybrid Target Attention architecture. Specifically, our model consists of stacked blocks; to mitigate the quadratic bottleneck, each block interleaves one full attention layer with efficient Target Attention layers. This hybrid design allows the model to capture global dependencies while significantly reducing the computational overhead compared to standard Transformers.
We first introduce the Full Attention Layer and the Target Attention Layer, respectively. In the Full Attention Layers, we first apply group layer normalization to the input tokens: H-tokens from different historical sequences are assigned to different groups; similarly, R-tokens from different real-time sequences are assigned to different groups; and T-tokens from different scenarios also belong to distinct groups. This design enables the model to better adapt to the diverse distributions arising from different sequences and token features:
Finally, the normalized token embeddings are fed into the HSTU for sequence modeling. We adopt Grouped-Query Attention (GQA) to further reduce computational overhead. Formally, a full attention layer at the -th layer of our hybrid attention architecture is defined as follows:
where is the number of query heads, is the number of key-value heads, is the number of query heads per key-value group; and denote linear layers; and are nonlinearity; and is our dynamic mask matrix to aviod information leakage. Because it was sorted in advance according to timestamps, we can construct a dynamic mask according to the following rules:
The H-tokens are visible to all tokens;
The R-tokens are visible only to tokens with later timestamps to avoid information leakage;
The T-tokens are visible only to themselves.
Figure 2 illustrates an example of our dynamic mask.
In the Target Attention Layers, we update only the embeddings of T-tokens, while the embeddings of other tokens from the previous layer are passed to the next layer via a shortcut connection. First, we extract the sub-matrix corresponding to the T-tokens from the normalized token embedding matrix and the dynamic mask matrix:
Then, we apply HSTU with GQA to update only the embeddings of T-tokens. Formally, a target attention layer at the -th layer of our hybrid attention architecture is defined as follows:
Finally, we concatenate the new embeddings of the T-tokens with the embeddings from the previous layer of other tokens to generate the output of the target attention layer.
The embeddings of the T-tokens from the final layer are input into an MMoE(Ma et al., 2018a) module to compute prediction scores for various objectives across multiple scenarios.
By employing this hybrid architecture, we reduce the computational complexity from to , where . Our experiments demonstrate that HTA achieves a 2x throughput speedup in training without compromising model accuracy.
3.3. Training and Deployment Optimization
3.3.1. Data Pipeline of MTFM
In this section, we delineate the data pipeline design of MTFM. Following the precedent set by MTGR, we utilize user-level sample compression to diminish the cardinality of the training set. We further extend this technique into a multi-scenario user-level compression framework, which is shown in Figure 3. Specifically, we decouple the feature space into two distinct components: scenario-agnostic (shared) features, such as , and scenario-specific features, such as . We first perform user-level aggregation on the scenario-specific features within their respective local scenarios. These aggregated components are then concatenated column-wise to construct a unified feature representation, which is finally merged with the scenario-agnostic data at the user level. This pipeline effectively circumvents computational redundancy and storage overhead, significantly improving systemic robustness and resource utilization efficiency.
3.3.2. Training Optimization
We conduct comprehensive optimizations to the offline training framework:
CPU-GPU Pipeline Optimization. In the domain of Large Language Models (LLMs), text undergoes tokenization with all features residing in GPU memory, enabling efficient end-to-end training. In contrast, recommendation systems face substantial feature engineering overhead, with numerous features stored in CPU memory. This architectural discrepancy creates blocking issues between GPU-resident models and CPU-resident features. The feature processing and model fusion pipelines involve extensive Host-Device data synchronization and serial dependencies, introducing multiple synchronization points where the Host must wait for Device-side operations to complete before proceeding. These CPU-GPU pipeline stalls severely degrade overall training performance.
To address this challenge, we systematically profiled the training framework using CUDA Profiler to identify CPU/GPU blocking bottlenecks. Our optimization strategy encompasses two key aspects: (1) eliminating synchronization points to enable CPU-GPU execution overlap, and (2) optimizing Device-side memory operations to reduce Device-to-Device (D2D) copy overhead. To implement the latter, we optimized frequent tensor indexing and assignment operations by consolidating multiple D2D copy operations into single atomic operations. Experimental results demonstrate that these systemic refinements improve training throughput by 20%.
Custom Kernel Development. First, we present an enhanced FlashAttention-2 (Dao, 2023) kernel designed to overcome the limitations of dynamic masking. While existing implementations are primarily optimized for standard causal masks, they often suffer from significant memory inefficiencies due to discontinuous access patterns when handling dynamic masks. To resolve this, we introduce contiguous and aligned mask layouts that facilitate high-throughput asynchronous copying. We also employ a fine-grained shared memory management strategy: we utilize available capacity to cache mask tiles during the forward pass, and implement temporal multiplexing of masks and intermediate tensors during the backward pass. This orchestration effectively overlaps mask-loading latency with the computation pipeline.
Furthermore, we implement fused operators for Group Layer Normalization (GLN) and dynamic mask construction using the Triton (Tillet et al., 2019). Native PyTorch implementations employ a Gather-Compute-Scatter pattern for GLN and fragmented operations for dynamic mask generation, resulting in excessive kernel launch overhead and inefficient memory access patterns. Leveraging Triton, we eliminate redundant intermediate memory transactions through operator fusion, enhance bandwidth utilization via vectorized memory access, and design group-wise parallel computation schemes to fully exploit GPU parallelism.
Through these kernel-level optimizations, we achieved an additional 57% throughput improvement.
3.3.3. Inference Optimization
To enhance the inference efficiency of the MTFM model in large-scale industrial scenarios, we propose a comprehensive optimization framework spanning the sparsity, kernel, and system levels:
Sparse Matrix Multiplication. Leveraging the Sparse Tensor Cores within the NVIDIA Ampere architecture, we implement 2:4 structured pruning on the linear projection layers (UVQK and Output) of the HSTU module. This technique achieves a 50% reduction in memory footprint and exploits double the theoretical peak throughput for matrix multiplication, significantly reducing latency without sacrificing predictive performance. This optimization yields approximately 10% throughput improvement and reduces inference latency by approximately 0.2ms.
Attention Computation Optimization. To address the irregular sparsity inherent in dynamic masks, we propose a fine-grained computation skipping strategy. This approach dynamically prunes redundant operations stemming from sequence padding and leverages domain-specific constraints to bypass invalid computation (e.g., zero-weight attention from user to target features). This optimization yields approximately 5% throughput improvement.
Scenario-aware Deployment and Other Optimizations. Since MTFM is trained across multiple scenarios but deployed on independent sub-scenarios, we decompose the global computational graph into scenario-specific subgraphs. This modularization eliminates redundant operations associated with unnecessary scenario-specific features, thereby optimizing resource utilization. Furthermore, we leverage BF16 half-precision inference and the M-Falcon (Zhai et al., 2024) algorithm to accelerate inference and maximize system throughput.
4. Experiments
In this section, we conduct extensive experiments on our real- world production datasets to validate the effectiveness of our proposed framework MTFM and answer the following research questions:
RQ1: How does the MTFM model perform across multiple heterogeneous multi-scenario tasks compared to other baseline models?
RQ2: How does the hybrid attention architecture in MTFM affect model training and inference performance?
RQ3: How does the MTFM model perform in terms of scalability, including model computational complexity scalability and data scalability?
RQ4:Can MTFM provide interpretable insights into cross-scenario token interactions?
RQ5: What is the online recommendation performance of MTFM for various heterogeneous scenarios in production environments?
4.1. Experiments Setup
4.1.1. Datasets
| Scenario | #Users | #Item | #Exposure | #Click | #Purchase |
| HP | 240M | 4.23M | 18.53B | 1.08B | 176.77M |
| PHF | 151M | 8.07M | 15.29B | 359.14M | 104.73M |
| SQS | 44M | 0.98M | 2.24B | 85.34M | 9.92M |
To accommodate the heterogeneous multi-scenario business data proposed in this work, we construct our offline experimental dataset from a real-world industrial food delivery recommendation system (Meituan) using user behavior logs. The dataset comprises three primary recommendation scenarios: Homepage (HP) Restaurant Recommendation, Shenqiangshou(SQS) Coupon-Package Recommendation , and Pinhaofan(PHF) Food Recommendation, with their respective data statistics summarized in Table 1. Specifically, HP serves as the highest-exposure scenario where restaurant cards are recommended to users for order placement, while the SQS and PHF scenarios are food-centric with distinct business objectives (detailed in 4.1.2). Over the past years, these scenarios have been continuously optimized both in business operations and recommendation architecture, making them the focal points for model innovation and industrial impact within Meituan’s ecosystem.
4.1.2. Metrics & Baselines
We evaluate model performance primarily using AUC (Area Under the Curve) and GAUC (Grouped AUC) to assess prediction quality across diverse scenarios. For HP and PHF, we focus on CTR and CTCVR to measure exposure and conversion effectiveness. In the SQS, we extend our evaluation to include package redemption efficiency via two specialized metrics: IMD (Immediate Redemption within 30 minutes) and WRITE (Redemption within 24 hours).
To ensure rigorous comparison, we categorize baseline models according to their architecture types: 1) General recommenders including DCNv2 (Wang et al., 2021), MMoE (Ma et al., 2018b), and RankMixer (Zhu et al., 2025), which are widely adopted for feature interaction and multi-task learning. 2) Generative Ranking Models, such as MTGR (Han et al., 2025) and OneTrans (Zhang et al., 2025a), representing the latest advances in generative sequence modeling. 3) Multi-Scenario Recommendation Models, such as STAR (Sheng et al., 2021b) and PEPNet (Chang et al., 2023), which integrate multiple scenarios into a unified framework, allowing adaptive scenario modeling. Notably, models in (1) and (2) are trained on a single-scenario dataset, whereas (3) and (4) leverage a multi-scenario dataset for training.
4.1.3. Hyper-parameter Settings.
For offline experiments, we use the Adam optimizer (learning rate ) with hyperparameters set to , and .
| Homepage Recommendation (HP) | Food Recommendation (PHF) | ||||||||
| Method | CTR | CTCVR | CTR | CTCVR | |||||
| AUC | GAUC | AUC | GAUC | AUC | GAUC | AUC | GAUC | ||
| DCNv2 | 0.7664 | 0.6853 | 0.8780 | 0.6451 | 0.7683 | 0.7236 | 0.8586 | 0.7555 | |
| MMoE | 0.7664 | 0.6857 | 0.8782 | 0.6454 | 0.7718 | 0.7282 | 0.8640 | 0.7597 | |
| General Recommenders | Rankmixer | 0.7665 | 0.6860 | 0.8789 | 0.6464 | 0.7711 | 0.7270 | 0.8628 | 0.7590 |
| OneTrans | 0.7672 | 0.6944 | 0.8774 | 0.6497 | 0.7832 | 0.7373 | 0.8827 | 0.7735 | |
| Generative Ranking Method | MTGR | 0.7679 | 0.6951 | 0.8776 | 0.6491 | 0.7883 | 0.7398 | 0.8879 | 0.7771 |
| STAR | 0.7669 | 0.6882 | 0.8780 | 0.6482 | 0.7821 | 0.7298 | 0.8688 | 0.7660 | |
| Multi-Scenario Method | PEPNet | 0.7672 | 0.6895 | 0.8790 | 0.6489 | 0.7866 | 0.7328 | 0.8721 | 0.7693 |
| Our Proposed Method | MTFM | 0.7689 | 0.6954 | 0.8806 | 0.6507 | 0.7940 | 0.7474 | 0.8892 | 0.7824 |
| Coupon-Package Recommendation (SQS) | |||||||||
| Method | CTR | CTCVR | IMD | WRITE | |||||
| AUC | GAUC | AUC | GAUC | AUC | GAUC | AUC | GAUC | ||
| DCNv2 | 0.8290 | 0.7789 | 0.9057 | 0.8227 | 0.9074 | 0.8254 | 0.9072 | 0.8222 | |
| MMoE | 0.8449 | 0.7842 | 0.9073 | 0.8243 | 0.9082 | 0.8252 | 0.9056 | 0.8230 | |
| General Recommenders | Rankmixer | 0.8472 | 0.7881 | 0.9094 | 0.8281 | 0.9105 | 0.8268 | 0.9080 | 0.8279 |
| OneTrans | 0.8454 | 0.7994 | 0.9091 | 0.8271 | 0.9089 | 0.8279 | 0.9055 | 0.8248 | |
| Generative Ranking Method | MTGR | 0.8454 | 0.7997 | 0.9097 | 0.8282 | 0.9095 | 0.8291 | 0.9059 | 0.8258 |
| STAR | 0.8470 | 0.7863 | 0.9079 | 0.8245 | 0.9081 | 0.8257 | 0.9055 | 0.8233 | |
| Multi-Scenario Method | PEPNet | 0.8511 | 0.7892 | 0.9081 | 0.8251 | 0.9089 | 0.8268 | 0.9066 | 0.8239 |
| Our Proposed Method | MTFM | 0.8624 | 0.8027 | 0.9119 | 0.8301 | 0.9117 | 0.8319 | 0.9079 | 0.8288 |
4.2. Overall Performance Comparison (RQ1)
Table 2 presents the main experimental results for three representative scenarios: HP, PHF, and SQS. Several key insights can be drawn. First, RankMixer delivers competitive results among general recommenders by effectively supporting model scaling; it marginally exceeds other baselines on HP while achieving substantial gains on SQS, outperforming MMoE by 0.38pp in CTCVR GAUC and 0.49pp in WRITE GAUC. However, in the PHF setting, its performance remains superior to DCNv2 but trails behind MMoE. Second, multi-scenario methods such as STAR and PEPNet generally outperform general baselines but fall short of generative ranking models like MTGR and OneTrans. Notably, these methods exhibit a pronounced ”see-saw effect”; for instance, PEPNet surpasses RankMixer in HP and PHF but underperforms in the SQS scenario. In contrast, generative models consistently demonstrate superior robustness, with MTGR slightly edging out OneTrans across most metrics. Ultimately, our proposed MTFM model achieves state-of-the-art results across nearly all scenarios and tasks. The consistent dominance of MTFM underscores its capability to not only mitigate the see-saw effect but, more importantly, to capitalize on the scaling law by efficiently leveraging extensive multi-scenario data to further enhance model’s performance.
4.3. Efficiency Analysis of HTA (RQ2)
| Configuration | Batch | AUC | GAUC | Throughput | Memory |
| (Target:Full) | Size | (samples/s) | (GB) | ||
| Full Attention Only | |||||
| (0:16)*1 | 1 | 0.7514 | 0.6818 | 390 | 66.97 |
| Target Attention Only | |||||
| (16:0)*1 | 1 | 0.7497 | 0.6806 | 575 | 32.64 |
| Hybrid Configurations | |||||
| (1:1)*8 | 1 | 0.7508 | 0.6820 | 497 | 38.00 |
| (3:1)*4 | 1 | 0.7506 | 0.6821 | 547 | 34.08 |
| (5:1)*3 | 1 | 0.7504 | 0.6811 | 535 | 37.52 |
| (3:1)*4 (w/o GQA) | 2 | 0.7506 | 0.6822 | 660 | 70.16 |
| (3:1)*4 (MTFM) | 2 | 0.7506 | 0.6822 | 780 | 67.49 |
In this section, we conducted extensive experiments to explore different configurations of the hybrid target attention. We performed experiments using another 7-day sample for training. We denote the various network architecture configurations as , where represents the total number of blocks. Within each block, and signify the number of Target Attention layers and Full Attention layers, respectively. We further evaluate a special case, Target Attention Only, where the model degenerates into the lazy decoder as seen in OneRecV2(Zhou et al., 2025b).
We evaluate the impact of varying the ratio between target and full attention layers (Table 3). The 1:1 and 3:1 configurations maintain or even slightly improve performance relative to full attention while delivering significant gains in both memory efficiency and training throughput. When the sparsity ratio is further increased to 5:1, performance begins to decline by 0.07pp, with the degradation widening to 0.12pp under a pure target attention setup. Additionally, the integration of GQA provides further gains in throughput alongside additional reductions in memory consumption. Finally, our HTA achieves a 2x speedup in training throughput compared to the full attention mechanism.
4.4. Scalability (RQ3)
In this section, we systematically evaluate the scalability of MTFM for multi-scenario recommendation (MSR) by examining its performance across varying model scales and training data volumes.
4.4.1. Scalability with Model Size in MSR
We evaluated its performance across various model scales (ranging from 10x to 70x), using MMoE as the competitive baseline. We employ CTCVR GAUC as the primary metric across three distinct scenarios: SQS, HP, and PHF. As illustrated in Figure 4(a), MTFM exhibits robust scaling laws with consistent slopes across all scenarios, underscoring its superior scalability regarding model capacity in MSR tasks.
4.4.2. Scalability with Training Data Volumes in MRS
As depicted in Figure 4(b), we investigate the impact of the data scale on MTFM by selecting models of three different sizes and observing their performance fluctuations in SQS’s CTCVR GAUC throughout the continual training phase. Empirical results reveal that across all model scales, performance improves consistently as the volume of training tokens increases. In experiments with larger data scales, we further observe that the performance gap between models of different sizes widens gradually as the number of training tokens increases. This robust scaling trend underscores MTFM’s data scalability and demonstrates its ability to consistently convert increased training volumes into performance improvements.
4.5. Interpretability Analysis (RQ4)
To understand how MTFM models scenario-specific characteristics and captures differentiated cross-scenario knowledge, we visualize the attention heatmaps for representative samples in Figure 5. The horizontal axis denotes T-Tokens across scenarios, while the vertical axis represents H-Tokens. Our observations include: 1) T-Tokens in all scenarios effectively aggregate information from H-Tokens across the entire multi-scenario space. 2) Distinct learning patterns emerge across scenarios; for instance, T-Tokens in HP and SQS exhibit higher attention weights toward their own H-Tokens compared to those of PHF, whereas PHF T-Tokens are more heavily indexed on their corresponding H-Tokens. Overall, the heatmaps reveal distinct block-wise attention patterns, demonstrating that MTFM is not only capable of learning from multi-scenario data but also possesses robust scenario-awareness.
4.6. Online Experiments (RQ5)
To validate MTFM’s real-world impact, we conducted A/B tests across two major online scenarios. The experimental traffic involved tens of millions of daily exposures, ensuring robust results. The baseline model was the current SOTA model, which has been continuously optimized and deployed for several years.
| Scenario | CTR | UV_CTCVR | ORDERS | LATENCY |
| SQS | +1.89% | +2.46% | +2.98% | -5ms |
| PHF | +1.53% | +1.03% | +1.45% | -6ms |
As shown in Table 4, MTFM delivers consistent performance improvements across all key business metrics. The observed online uplift in orders with MTFM is equivalent to the cumulative gains typically achieved over 2–3 rounds of model iteration in this business domain. In addition, we evaluate the latency of MTFM, where the results indicate that our system optimizations lead to better performance. These results clearly demonstrate that MTFM could deliver direct business value in real-world recommendation systems.
5. Conclusion
In this paper, we propose MTFM, a recommendation foundation model that leverages heterogeneous tokenization to achieve a truly alignment-free architecture across diverse domains. Through the co-design of algorithms, training, and deployment systems, MTFM simultaneously embodies the three quintessential properties: Scalability, Extensibility, and Efficiency. Extensive offline evaluations demonstrate significant performance gains, establishing a new scalable paradigm that successfully transcends the limitations of single-scenario silos. Finally, online A/B tests confirm MTFM’s real-world effectiveness, delivering substantial improvements in CTR and conversion volumes in Meituan.